Assume you have just been hired as a business manager of PizzaPalace, a regional pizza restaurant chain. The company’s EBIT was $50 million last year and is not expected to grow. The firm is currently financed with all equity, and it has 10 million shares outstanding. When you took your corporate finance course, your instructor stated that most firms’ owners would be financially better off if the firms used some debt. When you suggested this to your new boss, he encouraged you to pursue the idea. As a first step, assume that you obtained from the firm’s investment banker the following estimated costs of debt for the firm at different capital structures:

Percent Financed with Debt,

  0%

20  

  8.0%

30  

  8.5  

40  

10.0  

50  

12.0  

If the company were to recapitalize, then debt would be issued and the funds received would be used to repurchase stock. PizzaPalace is in the 40% state-plus-federal corporate tax bracket, its beta is 1.0, the risk-free rate is 6%, and the market risk premium is 6%.

1. Using the free cash flow valuation model, show the only avenues by which capital structure can affect value.

Using complete sentences and academic vocabulary, please answer question. Please don't copy from other sources. Thank you
Answer #1
Assuming that capital expenditure equals depreciation tax shield, FCF is equal to
EBT*(1-t).
1) FCF for all equity situation = 50*(1-40%) = $            30.00 million
WACC = Cost of equity = 6%+1*6% = 12.00%
Value of the firm = 30/12% = $         250.00 million
Value of one share = 250/10 = $            25.00
2) FCF with 20% debt = (50-50*20%*8%)*(1-40%) = $            29.52 million
[Debt = $250 million * 20% = $50 million]
Levered beta = Unlevered beta * (1+ (1-t) (Debt/Equity)) = 1*(1+0.60*0.2/0.8) = 1.15
Cost of levered equity = 6%+1.15*6% = 12.90%
Cost of debt = 8%*(1-40%) = 4.80%
WACC = 12.90%*80%+4.80%*20% = 11.28%
Value of the firm = 29.52/11.28% = $         261.70 million
Value of equity = 261.70-10.00 = $         251.70 million
Number of shares = 240/25 = 9.60 million
Value of one share = 251.70/9.6 = $            26.22
As a result of recapitalization with debt to the extent of 20%, the value of the firm
and the value of one share stands increased.